# Large Language Models are In-Context Molecule Learners

### v1

@@ 3.2.2 Sequence Reversal
$\mathcal{P}(x_{1},y_{1})\oplus\mathcal{P}(x_{1},y_{1})\oplus...\oplus\mathcal{P}(x_{n},y_{n}),$ (4)
@@ 3.3 In-Context Molecule Tuning
In contrast, ICMA first employs the Cross-modal Retrieval and Post-retrieval Re-ranking to obtain a subset $D_{(x,y)}\subset D$ containing $n$ similar examples $\{(x_{i},y_{i})\}$, where $1\leq i\leq n$, from the training set. Then, these similar examples could be wrapped into the context $C$. Notably, different from the previous ICL objective, ICMA is also motivated to learn the mapping $f_{i}:x_{i}\rightarrow y_{i}$ inside the context examples in an obvious manner. Finally, the context $C$ and the aggregation of mappings =$\mathcal{F}=\{f_{1},f_{1},...,f_{n}\}$ learned from context will altogether contribute to the learning of $f:x\rightarrow y$. Therefore, the objective of ICMA can be represented as:
$\mathcal{L}^{I\!C\!M\!A}(\theta)=\underset{(x,y)\in D}{\sum}-\log p_{\theta}(y|x,C,\mathcal{F}),$ (7)
where $\mathcal{L}^{I\!C\!M\!A}(\theta)$ denotes the overall loss function.
@@ 4 Experiments
In this section, we aim to evaluate the effectiveness of ICMA. Firstly, we introduce the experimental settings. Then, we compare ICMA with selected baseline models on ChEBI-20 dataset and further test ICMA on a smaller dataset, PubChem324k. Furthermore, we also study the factors that will affect the performance of IMCA, including different foundation models, retrieval algorithms, context settings, and model scales. Lastly, an ablation study is also conducted to justify the design of Post-retrieval Re-ranking components.
@@ 4.1 Experimental Settings
We will first detail our experiment settings. All the hyper-parameters are illustrated in Appendix A Table 10. If not specifically stated, the cutoff length is set to 1024 and n_shot is set to 2 to control the parameters. Notably, for LLMs with over 7 billion parameters, we apply LoRA  Hu et al. (2021) to save the GPU memory and accelerate computation. Otherwise, we fine-tune the full model of LLMs. For the dataset, we apply two different molecule-caption translation datasets, ChEBI-20  Edwards et al. (2022) and PubChem324k  Liu et al. (2023b). The details of the datasets are shown in Table 1.
@@ 4.2 Performance Comparison
ICMA(Mistral-7B)${}_{2,2048}$ 0.651 0.581 0.686 0.550 0.625 0.661
Table 2: Mol2Cap results on ChEBI-20 dataset (Best, Second Best).
@@
ICMA(Mistral-7B)${}_{4,2048}$ 0.855 0.460 18.73 0.916 0.837 0.789 0.958
Table 3: Cap2Mol results on ChEBI-20 dataset (Best, Second Best).
@@
Mol2Cap. As illustrated in Table 2, Galactica-125M with SFT has already shown competitive performance to previous baselines due to its pre-training on scientific corpora. However, ICMA could still improve the performance of Galactica-125M by 12.8% and 8.3% considering the BLEU-4 and ROUGE-L scores on ChEBI-20 dataset. With only 125 million parameters, ICMA(Galactica-125M) can beat MolT5-large, which owns 780 million parameters. Meanwhile, the general LLM, Mistral-7B with naive SFT, only achieves a performance that is slightly better than MolT5-base, despite the fact that Mistral-7B is 70 times larger than MolT5-base. This outcome is not surprising because Mistral-7B is not specifically designed or pre-trained for biomolecular purposes. It also reveals that although general LLMs have illustrated powerful capabilities with billions of parameters, few works have adapted them to the biomolecular domain due to their unsatisfactory fine-tuning performance. However, with ICMA, Mistral-7B could easily demonstrate its superior in-context learning and reasoning capabilities and perform its advantage of parameters. As a result, ICMA(Mistral-7B) achieves the best performance across all the models, obatining 0.581 BLEU-4 and 0.661 METEOR scores on ChEBI-20 dataset, which is even better than domain-specific pre-trained Galactica-125M.
@@
ICMA(Mistral-7B)2,2048ß${}_{2,2048ß}$ 0.386 0.314 0.483 0.345 0.431 0.443
Table 4: Mol2Cap results on PubChem324k dataset (Best, Second Best).
@@
ICMA(Mistral-7B)${}_{1,2048}$ 0.450 0.065 77.01 0.764 0.624 0.504 0.891
Table 5: Cap2Mol results on PubChem324k dataset (Best, Second Best).
@@ 4.3 Study of Retrieval Algorithms
Table 6: Performance comparison of different retrieval algorithms for Mol2Cap task (Best, Second Best).
@@
Table 7: Performance comparison of different retrieval algorithms for Cap2Mol task (Best, Second Best).
We study the influence of retrieval algorithms to illustrate the importance of retrieval qualities. For molecule retrieval, we compare the new proposed Molecule Graph Retrieval using Mole-BERT with random retrieval and the Morgan FTS retrieval. As shown in Table 6, Molecule Graph Retrieval using Mole-BERT achieves the best results on all of the metrics, proving its superiority in molecule retrieval. For the caption retrieval, we compare BM25 Caption Retrieval with random retrieval and SBERT retrieval under the framework of ICMA. As depicted in Table 7, BM25 Caption Retrieval illustrates its excellent performance among the three caption retrieval methods. Notably, SBERT retrieval achieves worse score than random retrieval because it losses detailed information during the compression, further proving the importance of details in the caption retrieval.
@@ 4.4 Study of Context Settings
Figure 3: The model performance with the change of context settings, including number of examples (i.e., n_shot) and cutoff length. Mol2Cap Results (Left) and Cap2Mol Results (Right).
In ICMA, the context settings, including the example number and cutoff length, are also important to its performance. The increasing of context example number will also drastically add the context length. However, if the context length is larger than the cutoff length, then the training will be insufficient because most of the inputs are cropped and the information is lost. In this case, we want to make sure that most of the context examples fit in the cutoff length. Considering that the input length limitation of Galactica-125M is 2048, we test the cutoff length within the range of {512, 1024, 1536, 2048} and the example number from 1 to 4 for analysis. As illustrated in Figure 3, when the example number increases, the performance would become worse if the cutoff length is too short. For example, when the cutoff length is set to 512, the context is not complete for LLMs to consistently learn valuable knowledge from it, leading to worse results. However, when the cutoff length is long enough, the performance mainly becomes better as the increased example number provides more information to help LLMs make accurate predictions.
@@ 4.5 Study of Scaling Law
As illustrated in Figure 4, with the increase of model scale, the performance of ICMA also improves. Especially in Mol2Cap task, the METEOR score has been drastically improved when the model parameters scale from 1.3 billion to 7 billion. The improvement of molecule understanding significantly benefits from the language generation capability as the model scale reaches a certain level. Meanwhile, for Cap2Mol task, the accuracy for predicting exactly matched molecules demonstrates a slightly different pattern to Mol2Cap task. The model’s performance in generating exact matched molecules seems to have a log-linear relationship with the model scales, showing that larger language models could explore the molecular space better owing to their more powerful capabilities. Overall, ICMA allows general LLMs with billion parameters to achieve similar or better performance to these domain-specific pre-trained LLMs in the molecule-caption translation task, showing that LLMs are inherently in-context molecule learners.
@@ 4.6 Ablation Study
Table 8: Ablating components of Post-retrieval Re-ranking for Mol2Cap task (Best, Second Best).
@@
Table 9: Ablating components of Post-retrieval Re-ranking for Cap2Mol task (Best, Second Best).
@@ A.2 Context Templates
A.3 Case Study

### v2

@@ 3.2.2 Sequence Reversal
$\mathcal{P}(x_{1},y_{1})\oplus\mathcal{P}(x_{2},y_{2})\oplus...\oplus\mathcal{P}(x_{n},y_{n}),$ (4)
@@ 3.3 In-Context Molecule Tuning
In contrast, ICMA does not simply conduct the next token prediction on the output part but learns the entire input in an auto-regressive manner. ICMA first employs the Hybrid Context Retrieval and Post-retrieval Re-ranking to obtain a subset $D_{(x,y)}\subset D$ containing $n$ similar examples $\{(x_{i},y_{i})|1\leq i\leq n\}$, from the training set. Different from the previous ICL objective, ICMA is also motivated to learn the mapping $f_{i}:x_{i}\rightarrow y_{i}$ inside the context examples in an obvious manner. Notably, molecule-caption mapping is the most informative part in the context examples because similar molecule sub-structures inherit similar characteristics. For example, RCOOH (carboxyl group) usually indicates that the molecule is an acid. In this way, learning the alignment between functional groups and molecule captions in the context examples could benefit the final prediction. For simplicity, we could assume that context examples are independent of each other as the molecule-caption mapping plays the most important role in the prediction. In this case, the aggregation of mappings $\mathcal{F}_{(x,y)}=\{f_{1},f_{2},...,f_{n}\}$ could be learned from context and will altogether contribute to the final prediction with the corresponding context $C_{(x,y)}$, which wraps the context examples into the input text. Therefore, the objective of ICMA can be represented as:
$\mathcal{L}(\mathcal{F}_{(x,y)})=\underset{(x_{i},y_{i})\in D_{(x,y)}}{\sum}-\log p_{\theta}(y_{i}|x_{i}),$ (7)
$\mathcal{L}^{I\!C\!M\!A}(\theta)\!=\!\underset{(x,y)\in D}{\sum}\!\left(\!-\log p_{\theta}(y|x,C_{(x,y)})\!+\!\mathcal{L}(\mathcal{F}_{(x,y)})\right)\!,$ (8)
where $\mathcal{L}(\mathcal{F}_{(x,y)})$ represents the aggregated mapping loss for molecule-caption pair $(x,y)$, while $\mathcal{L}^{I\!C\!M\!A}(\theta)$ denotes the overall loss function.
@@ 4 Experiments
In this section, we aim to evaluate the effectiveness of ICMA. Firstly, we introduce the experimental settings. Then, we compare ICMA with selected baseline models on the ChEBI-20 dataset and further test ICMA on a smaller dataset, PubChem324k. Furthermore, we also study the factors that will affect the performance of IMCA, including different foundation models, retrieval algorithms, context settings, and model scales. Lastly, an ablation study is also conducted to justify the design of Post-retrieval Re-ranking components.
@@ 4.1 Experimental Settings
We will first detail our experiment settings. All the hyper-parameters are illustrated in Appendix A Table 10. If not specifically stated, the cutoff length is set to 1024 and n_shot is set to 2 to control variable. Notably, for LLMs with over 7 billion parameters, we apply LoRA  Hu et al. (2021) to save the GPU memory and accelerate computation. Otherwise, we fine-tune the full model of LLMs. For the dataset, we apply two different molecule-caption translation datasets, ChEBI-20  Edwards et al. (2021) and PubChem324k  Liu et al. (2023b). The details of the datasets are shown in Table 1.
@@ 4.2 Performance Comparison
ICMA(Mistral-7B)2,2048 0.651 0.581 0.686 0.550 0.625 0.661
Table 2: Mol2Cap results on ChEBI-20 dataset (Best, Second Best). Here, the results of MolT5-base and MolT5-large are domain-specific pre-training & fine-tuning results retrieved from Edwards et al. (2022), while MolReGPT(GPT-3.5-turbo) and MolReGPT(GPT-4-0314) are prompting & in-context learning results retrieved from Li et al. (2023). More importantly, Galactica-125M and Mistral-7B demonstrate vanilla supervised fine-tuned results, while ICMA(Galactica-125M) and ICMA(Mistral-7B) illustrate the ICMA results.
@@
ICMA(Mistral-7B)4,2048 0.855 0.460 18.73 0.916 0.837 0.789 0.958
Table 3: Cap2Mol results on ChEBI-20 dataset (Best, Second Best). Here, the results of MolT5-base and MolT5-large are domain-specific pre-training & fine-tuning results retrieved from Edwards et al. (2022), while MolReGPT(GPT-3.5-turbo) and MolReGPT(GPT-4-0314) are prompting & in-context learning results retrieved from Li et al. (2023). More importantly, Galactica-125M and Mistral-7B demonstrate vanilla supervised fine-tuned results, while ICMA(Galactica-125M) and ICMA(Mistral-7B) illustrate the ICMA results.
@@
Mol2Cap. As illustrated in Table 2, Galactica-125M with SFT has already shown competitive performance to previous baselines due to its pre-training on scientific corpora. However, ICMA could still improve the performance of Galactica-125M by 12.8% and 8.3% considering the BLEU-4 and ROUGE-L scores on ChEBI-20 dataset. With only 125 million parameters, ICMA(Galactica-125M) can beat MolT5-large, which owns 780 million parameters. Meanwhile, the general LLM, Mistral-7B with naive SFT, only achieves a performance that is slightly better than MolT5-base, despite the fact that Mistral-7B is 70 times larger than MolT5-base. This outcome is not surprising because Mistral-7B is not specifically designed or pre-trained for biomolecular purposes. It also reveals that although general LLMs have illustrated powerful capabilities with billions of parameters, few works have adapted them to the biomolecular domain due to their unsatisfactory fine-tuning performance. However, with ICMA, Mistral-7B could easily demonstrate its superior in-context learning and reasoning capabilities and perform its advantage of parameters. As a result, ICMA(Mistral-7B) achieves the best performance across all the models, obtaining 0.581 BLEU-4 and 0.661 METEOR scores on the ChEBI-20 dataset, which is even better than the domain-specific pre-trained Galactica-125M.
@@
ICMA(Mistral-7B)2,2048ß 0.386 0.314 0.483 0.345 0.431 0.443
Table 4: Mol2Cap results on PubChem324k dataset (Best, Second Best). Here, Galactica-125M and Mistral-7B demonstrate vanilla supervised fine-tuned results, while ICMA(Galactica-125M) and ICMA(Mistral-7B) illustrate the ICMA results.
@@
ICMA(Mistral-7B)1,2048 0.450 0.065 77.01 0.764 0.624 0.504 0.891
Table 5: Cap2Mol results on PubChem324k dataset (Best, Second Best). Here, Galactica-125M and Mistral-7B demonstrate vanilla supervised fine-tuned results, while ICMA(Galactica-125M) and ICMA(Mistral-7B) illustrate the ICMA results.
@@ 4.3 Study of Retrieval Algorithms
Table 6: Performance comparison of different retrieval algorithms for Mol2Cap task (Best, Second Best). The backbone is ICMA(Galactica-125M)2,1024.
@@
Table 7: Performance comparison of different retrieval algorithms for Cap2Mol task (Best, Second Best). The backbone is ICMA(Galactica-125M)2,1024.
We study the influence of retrieval algorithms to illustrate the importance of retrieval qualities. For molecule retrieval, we compare the new proposed Molecule Graph Retrieval using Mole-BERT with random retrieval and the Morgan FTS retrieval. As shown in Table 6, Molecule Graph Retrieval using Mole-BERT achieves the best results on all of the metrics, proving its superiority in molecule retrieval. For the caption retrieval, we compare BM25 Caption Retrieval with random retrieval and SBERT retrieval under the framework of ICMA. As depicted in Table 7, BM25 Caption Retrieval illustrates its excellent performance among the three caption retrieval methods. Notably, SBERT retrieval achieves a worse score than random retrieval because it loses detailed information during the compression, further proving the importance of details in caption retrieval.
@@ 4.4 Study of Context Settings
Figure 3: The model performance with the change of context settings, including the number of examples (i.e., n_shot) and cutoff length. Mol2Cap Results (Left) and Cap2Mol Results (Right).
In ICMA, the context settings, including the example number and cutoff length, are also important to its performance. The increase of context example number will also drastically add to the context length. However, if the context length is larger than the cutoff length, then the training will be insufficient because most of the inputs are cropped, and the information is lost. In this case, we want to make sure that most of the context examples fit in the cutoff length. Considering that the input length limitation of Galactica-125M is 2048, we test the cutoff length within the range of {512, 1024, 1536, 2048} and the example number from 1 to 4 for analysis. As illustrated in Figure 3, when the example number increases, the performance becomes worse if the cutoff length is too short. For example, when the cutoff length is set to 512, the context is not complete for LLMs to consistently learn valuable knowledge from it, leading to worse results. However, when the cutoff length is long enough, the performance mainly becomes better as the increased example number provides more information to help LLMs make accurate predictions.
@@ 4.5 Study of Scaling Law
As illustrated in Figure 4, with the increase of model scale, the performance of ICMA also improves. Especially in Mol2Cap task, the METEOR score has been drastically improved when the model parameters scale from 1.3 billion to 7 billion. The improvement of molecule understanding significantly benefits from the language generation capability as the model scale reaches a certain level. Meanwhile, for Cap2Mol task, the accuracy for predicting exactly matched molecules demonstrates a slightly different pattern from the Mol2Cap task. The model’s performance in generating exact-matched molecules seems to have a log-linear relationship with the model scales, showing that larger language models could explore the molecular space better owing to their more powerful capabilities. Overall, ICMA allows general LLMs with billion parameters to achieve similar or better performance to these domain-specific pre-trained LLMs in the molecule-caption translation task, showing that LLMs are inherently in-context molecule learners.
@@ 4.6 Ablation Study
Table 8: Ablating components of Post-retrieval Re-ranking for Mol2Cap task (Best, Second Best). The backbone is ICMA(Galactica-125M)2,1024.
@@
Table 9: Ablating components of Post-retrieval Re-ranking for Cap2Mol task (Best, Second Best). The backbone is ICMA(Galactica-125M)2,1024.
@@ A.2 Context Templates
A.3 Further Discussion
In this section, we conduct some further experiments to provide extra insights and support our method. First, we test and compare the performances of different backbone LLMs with and without ICMA. Then, we compare ICMA with some other baseline models.
A.3.1 Results with More Backbone LLMs
To highlight the model agnosticism of ICMA, we have expanded our experiments on two more backbone LLMs, Galactica-1.3B and Llama-2-7b-chat-hf. The results are shown in Tables 11 and 12. It is evident that ICMA enhances the performance of both models, which further proves the versatility of ICMA.
[…]
Galactica-1.3B 0.589 0.507 0.638 0.484 0.574 0.597
ICMA(Galactica-1.3B) 0.602 0.534 0.668 0.530 0.607 0.649
Llama2-7B 0.515 0.423 0.583 0.412 0.518 0.528
ICMA(Llama2-7B) 0.629 0.553 0.663 0.518 0.601 0.635
Table 11: Performance comparison of more backbone LLMs for Mol2Cap task on ChEBI-20 dataset. Here, we select Galactica-1.3B and Llama-2-7b-chat-hf for demonstration.
[…]
Galactica-1.3B 0.812 0.292 22.47 0.882 0.777 0.709 0.950
ICMA(Galactica-1.3B) 0.839 0.396 21.61 0.910 0.829 0.770 0.946
Llama2-7B 0.663 0.130 37.24 0.801 0.649 0.572 0.861
ICMA(Llama2-7B) 0.783 0.319 27.74 0.888 0.792 0.727 0.923
Table 12: Performance comparison of more backbone LLMs for Cap2Mol task on ChEBI-20 dataset. Here, we select Galactica-1.3B and Llama-2-7b-chat-hf for demonstration.
A.3.2 Comparison with Other Baseline Models
In our pursuit to comprehensively evaluate the efficacy of ICMA, we have extended our comparative analysis to include additional baseline models. Specifically, we first compare ICMA(Mistral-7B) with BioT5  Pei et al. (2023),  Liu et al. (2023a), and MolCA  Liu et al. (2023b) in the Mol2Cap task. The results, as detailed in Table 13, indicate that ICMA outperforms all the baseline models in terms of BLEU-2, BLEU-4, and METEOR scores, while also achieving the second-highest ROUGE scores, which further prove that ICMA could empower Mistral-7B to achieve SOTA or comparable performance.
Notably, compared to MolCA, a multi-modal method that leverages 2D molecular graphs to augment LLMs for the Mol2Cap task, our ICMA is capable of refining the alignment between molecular and textual representations not only for Mol2Cap task but also for Cap2Mol task, which is more important to the field of molecule discovery.
Furthermore, for the Cap2Mol task, we have included BioT5  Pei et al. (2023) and MolXPT  Liu et al. (2023a) in our comparative analysis. As presented in Table 14, ICMA achieves the highest scores in exact match and molecule fingerprints metrics, indicative of superior quality in molecule generation. However, it is also important to note that since the Mistral-7B model was not pre-trained on a large-scale chemical corpus, the validity of the generation is slightly below the two baseline models due to a deficiency in molecular knowledge.
[…]
BioT5  Pei et al. (2023) 0.635 0.556 0.692 0.559 0.633 0.656
MolXPT  Liu et al. (2023a) 0.594 0.505 0.660 0.511 0.597 0.626
MolCA  Liu et al. (2023b) 0.620 0.531 0.681 0.537 0.618 0.651
ICMA(Mistral-7B) 0.651 0.581 0.686 0.550 0.625 0.661
Table 13: Performance comparison with more baseline models for Mol2Cap task on ChEBI-20 dataset. Here, we select BioT5  Pei et al. (2023), MolXPT  Liu et al. (2023a), and MolCA  Liu et al. (2023b) for comparison.
[…]
BioT5  Pei et al. (2023) 0.867 0.413 15.10 0.886 0.801 0.734 1.000
MolXPT  Liu et al. (2023a) - 0.215 22.47 0.859 0.757 0.667 0.983
ICMA(Mistral-7B) 0.855 0.460 18.73 0.916 0.837 0.789 0.958
Table 14: Performance comparison of more baseline models for Cap2Mol task on ChEBI-20 dataset. Here, we include the results from BioT5  Pei et al. (2023) and MolXPT  Liu et al. (2023a) for comparison.
A.4 Case Study